### v2

@@ 6.1.3 Dimension-independent block acceptance rate
To test the dimension independent block acceptance rate in Proposition 4.1, we partition the original $512\times 512$ image into 4 sections of sizes $128\times 128$, $256\times 256$, $384\times 384$, and $512\times 512$. Furthermore, each section is partitioned into blocks of equal size $64\times 64$. Thus, the number of blocks in the sections of sizes $128\times 128$, $256\times 256$, $384\times 384$, and $512\times 512$ are $4$, $16$, $36$, and $64$, respectively. The 4 deblurring problems are shown on the left in Fig. 3.
We run the local & parallel MLwG given in Algorithm 2 with a step size of $\tau=7.44\times 10^{-6}$ on the 4 deblurring problems with different sizes. The step size is taken from a pilot run on the $512\times 512$ problem by targeting an acceptance rate of $0.547$ in each block and then taking the average of all block step sizes. For all problem sizes, we use a burn-in period of $31,250$ samples. We plot the acceptance rate for each block in Fig. 5, and see that the block acceptance rates are indeed dimension-independent.
@@ 6.1.4 Comparison to MALA
In this subsection, we compare the performance of the our method with MALA. For MALA, we use a diminishing step size adaptation during burn-in and target a step size of $0.547$. Further, the numbers of burn-in samples are listed in Table 2 and are chosen such that they increase linearly with the problem size. For MLwG, we use the same setting as in the previous tests.
We compare the sampling performance of MALA and MLwG in Table 2. MLwG yields in general much larger nESS than MALA, because it allows for a larger step size. Furthermore, the nESS of MLwG becomes even larger as the problem size increases. We attribute this to the diminishing constraining effect of the boundary condition associated with the convolution operator on the inner blocks as the dimension increases. In addition, it must be noted that MALA may not converge for the problem sizes $384\times 384$ and $512\times 512$ since the corresponding $\mathrm{max}~{}\mathrm{PSRF}>1.1$.
Table 2: Comparison of the local & parallel MLwG and MALA. The printed nESS is the mean over all pixels. For MLwG, the printed acceptance rate and step size are the mean over all blocks. The printed PSRF is the maximum over all pixels.
@@
nESS [%] MLwG 22.1 24.4 26.2 27.0
MALA 13.7 9.2 7.4 6.3
@@
Notice that the results from Table 2 also validate the dimension-independent convergence rate in Proposition 4.2 of MLwG. This is because MLwG produces for all problem sizes and with the same burn-in converged chains with roughly constant PSRF. In contrast, MALA requires significantly more burn-in with increasing dimension. We note that despite the clear differences in sampling performance between MLwG and MALA, we do not see clear differences in the sample mean or the widths of the $90\%$ CIs. Therefore, we do not include these results in this paper.
@@
We show the computing times in terms of samples per second in Fig. 6 and observe that the wall-clock time of MLwG remains almost constant and does not increase with the problem dimension. This is because the main computational effort of updating the $64\times 64$ blocks on each core remains constant and only more time is required for handling the increasing number of cores by the main process. For small problem sizes, the wall-clock time of MLwG is longer than that of MALA, because of the overhead of the parallelized implementation and the additional convolutions of fixed pixels in the local block likelihoods. However, since several updates are run in parallel in MLwG, its wall-clock time is eventually shorter than that of MALA, see the time for problem size $512\times 512$. Note that the total wall-clock time of MALA is actually significantly larger since it requires much more burn-in. The benefits of MLwG obviously come at the cost of CPU time, which increases linearly with the number of cores.
@@ 6.2 House
In this section, we use another test image with a different blurring kernel to compare our local & parallel MLwG with MALA.
@@ 6.2.1 Problem description
The degraded image is synthetically obtained via the observation model Eq. 3, where $\mathbf{A}\in\mathbb{R}^{d\times d}$ is obtained through the discrete PSF of a motion blur kernel with the length $17$ and the angle $45^{\circ}$. The noise is a realization of $\mathbf{\epsilon}\sim\mathcal{N}(\mathbf{0},0.01^{2}\mathbf{I})$. The true and the degraded image of “house” with the size $512\times 512$ are shown in the first row of Fig. 7.
Figure 7: The true image, the degraded image, the posterior sample mean and the width of the 90% sample CI.
We use the adaptive total variation approach in [23] to determine the rate parameter $\delta$ in the TV prior Eq. 7, and obtain $\delta=34.83$.
@@ 6.2.2 Posterior sampling via MLwG and MALA
Again, we apply both the local & parallel MLwG and MALA to sample the smoothed posterior defined in Eq. 14. In both methods, we target an acceptance rate of $0.547$, and in MLwG we adapt the step size individually in each block. In the second row of Figure 7, we show the posterior mean and the width of the 90% CI obtained via the samples from MLwG. The restoration results from MALA are neglected, since they are visually identical to the ones from MLwG. However, the data about the sample chains that we show in Table 3 reveals once again the superior performance of MLwG over MALA. Similar as observed in the previous test, MLwG allows for a significantly larger step size which leads to faster convergence and less correlated samples in terms of nESS.
Table 3: Sampling data of MLwG and MALA for the “house” image. The shown nESS is the mean over all samples. The shown step size and acceptance rate are the mean over all blocks. The maximum and median PSRF are with respect to all pixels.
nESS [%] $\bm{\tau}$ [10-6] $\bm{\alpha}$ [%] burn-in [103] max PSRF
MLwG 32.2 7.4 54.3 31.250 1.02
MALA 7.8 1.4 55.9 2000.000 1.09
@@ 7 Conclusions
Uncertainty quantification in imaging problems is usually a difficult task due to the high dimensionality of images. For image deblurring with TV prior, we present a dimension-independent MLwG sampling algorithm. By exploiting the sparse conditional structure of the posterior, the proposed algorithm has dimension-independent block acceptance rates and convergence rate, which are both theoretically proven and numerically validated. To enable the use of the MALA proposal density, we use a smooth approximation of the TV prior, and show that the introduced error is uniformly distributed over the pixels, and thus is dimension-independent. Moreover, through numerical studies, we find that the smoothed posterior converges quickly to the exact posterior and yields feasible results for uncertainty quantification.
@@ Acknowledgments
The work of RF and YD has been supported by a Villum Investigator grant (no. 25893) from The Villum Foundation. The work of SL has been partially funded by Singapore MOE grant A-8000459-00-00. The work of XT has been funded by Singapore MOE grants A-0004263-00-00 and A-8000459-00-00.
@@ A.1 Dimension-independent approximation error
By Kantorovich duality, the Wasserstein-$1$ distance of $\pi_{i}$ and $\pi_{\varepsilon,i}$ can be written as
$W_{1}(\pi_{i},\pi_{\varepsilon,i})=\sup_{\phi_{0}\in{\rm Lip}_{1}(\mathbb{R}^{q})}\int\phi_{0}(\mathbf{x}_{i})\left(\pi_{i}(\mathbf{x}_{i})-\pi_{\varepsilon,i}(\mathbf{x}_{i})\right)\mathrm{d}\mathbf{x}_{i},$
@@
$\phi(\mathbf{x}_{i})=\phi_{0}(\mathbf{x}_{i})-\int\phi_{0}(\mathbf{x}_{i})\pi_{\varepsilon,i}(\mathbf{x}_{i})\mathrm{d}\mathbf{x}_{i}=\phi_{0}(\mathbf{x}_{i})-\int\phi_{0}(\mathbf{x}_{i})\pi_{\varepsilon}(\mathbf{x})\mathrm{d}\mathbf{x}.$
Then $\phi\in{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$, the class of $1$-Lipschitz functions that are mean-zero w.r.t. $\pi_{\varepsilon,i}$:
(30) ${\rm Lip}_{1}^{0}(\pi_{\varepsilon,i}):=\left\{\phi\in{\rm Lip}_{1}(\mathbb{R}^{q}):\int\phi(\mathbf{x}_{i})\pi_{\varepsilon,i}(\mathbf{x}_{i})\mathrm{d}\mathbf{x}_{i}=0\right\}.$
Then by definition,
$\int\phi_{0}(\mathbf{x}_{i})\left(\pi_{i}(\mathbf{x}_{i})-\pi_{\varepsilon,i}(\mathbf{x}_{i})\right)\mathrm{d}\mathbf{x}_{i}=~{}$ $\int\phi(\mathbf{x}_{i})\left(\pi_{i}(\mathbf{x}_{i})-\pi_{\varepsilon,i}(\mathbf{x}_{i})\right)\mathrm{d}\mathbf{x}_{i}$
@@
Given $\phi(\mathbf{x}_{i})$, consider the Poisson equation for $u(\mathbf{x})$:
(31) $\Delta u(\mathbf{x})+\nabla\log\pi_{\varepsilon}(\mathbf{x})\cdot\nabla u(\mathbf{x})=\phi(\mathbf{x}_{i}).$
@@
$\int\phi(\mathbf{x}_{i})\pi(\mathbf{x})\mathrm{d}\mathbf{x}=~{}$ $\int\left(-\nabla u(\mathbf{x})\cdot\nabla\log\pi(\mathbf{x})+\nabla\log\pi_{\varepsilon}(\mathbf{x})\cdot\nabla u(\mathbf{x})\right)\pi(\mathbf{x})\mathrm{d}\mathbf{x}$
@@
Combine the above two inequalities, we have
(32) $\begin{split}\max_{i}W_{1}(\pi_{i},\pi_{\varepsilon,i})=~{}&\max_{i}\sup_{\phi\in{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})}\int\phi(\mathbf{x}_{i})\pi(\mathbf{x})\mathrm{d}\mathbf{x}\\ \leq~{}&2\lambda^{-1}c^{-1}\max_{j}\left\|\nabla_{j}\log\pi_{\varepsilon}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)}.\end{split}$
Note here $\nabla_{j}\log\pi$ is not defined pointwise, but it suffices to require that $\nabla_{j}\log\pi\in L^{1}(\pi)$. It suffices to control the right hand side of Eq. 32. By definition,
$\nabla_{j}\log\pi(\mathbf{x})-\nabla_{j}\log\pi_{\varepsilon}(\mathbf{x})=\nabla_{j}\varphi_{\varepsilon}(\mathbf{x})-\nabla_{j}\varphi_{0}(\mathbf{x})$
@@
$\qquad\cdot\left(\left(\mathbf{D}_{\alpha}^{(v)}\mathbf{x}\right)\left(\mathbf{D}_{\alpha,j}^{(v)}\right)^{\rm T}+\left(\mathbf{D}_{\alpha}^{(h)}\mathbf{x}\right)\left(\mathbf{D}_{\alpha,j}^{(h)}\right)^{\rm T}\right)=:\delta\sum_{\alpha=1}^{d}\mathbf{v}_{\alpha}(\mathbf{x}).$
Here we denote $\mathbf{D}_{\alpha,j}^{(v)}\in\mathbb{R}^{1\times q}$ as the $j$-th block of $\mathbf{D}_{\alpha}\in\mathbb{R}^{1\times d}$, and similar for $\mathbf{D}_{\alpha,j}^{(h)}$. To control $\left\|\mathbf{v}_{\alpha}(\mathbf{x})\right\|_{2}$, notice
$\left\|\left(\mathbf{D}_{\alpha}^{(v)}\mathbf{x}\right)\left(\mathbf{D}_{\alpha,j}^{(v)}\right)^{\rm T}+\left(\mathbf{D}_{\alpha}^{(h)}\mathbf{x}\right)\left(\mathbf{D}_{\alpha,j}^{(h)}\right)^{\rm T}\right\|_{2}\leq|\mathbf{D}_{\alpha}^{(v)}\mathbf{x}|\|\mathbf{D}_{\alpha,j}^{(v)}\|_{2}+|\mathbf{D}_{\alpha}^{(h)}\mathbf{x}|\|\mathbf{D}_{\alpha,j}^{(h)}\|_{2}$
$\leq\sqrt{2}\left(|\mathbf{D}_{\alpha}^{(v)}\mathbf{x}|+|\mathbf{D}_{\alpha}^{(h)}\mathbf{x}|\right)\leq 2\left((\mathbf{D}_{\alpha}^{(v)}\mathbf{x})^{2}+(\mathbf{D}_{\alpha}^{(h)}\mathbf{x})^{2}\right)^{1/2}=:2\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}$
Here we denote $\mathbf{D}_{\alpha}\mathbf{x}=(\mathbf{D}_{\alpha}^{(v)}\mathbf{x},\mathbf{D}_{\alpha}^{(h)}\mathbf{x})\in\mathbb{R}^{2}$. Notice $\mathbf{D}_{\alpha,j}^{(v)}=\mathbf{D}_{\alpha,j}^{(h)}=0$ if $\alpha$ does not live in block $j$ or a neighbor of block $j$, so that there are at most $(m+2)^{2}$ indices $\alpha$ s.t. $\mathbf{v}_{\alpha}(\mathbf{x})$ is nonzero. Therefore,
(33) $\begin{split}\|\nabla_{j}\log\pi(\mathbf{x})-\nabla_{j}\log&\pi_{\varepsilon}(\mathbf{x})\|_{L^{1}(\pi)}\leq\delta\cdot\mathbb{E}_{\mathbf{x}\sim\pi}\sum_{\alpha:\mathbf{v}_{\alpha}(\mathbf{x})\neq 0}\left\|\mathbf{v}_{\alpha}(\mathbf{x})\right\|_{2}\\ \leq~{}&2\delta(m+2)^{2}\max_{\alpha}\mathbb{E}_{\mathbf{x}\sim\pi}\left[1-\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}}{\sqrt{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{2}+\varepsilon}}\right].\end{split}$
Denote the function $I_{\alpha}(t):=\mathbb{E}_{\mathbf{x}\sim\pi}\left[1-\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}}{\sqrt{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{2}+t^{2}}}\right]$. Then $I_{\alpha}(0)=0$ and
$I_{\alpha}^{\prime}(t)=~{}$ $\mathbb{E}_{\mathbf{x}\sim\pi}\left[\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}t}{\left(\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{2}+t^{2}\right)^{3/2}}\right]\leq\mathbb{E}_{\mathbf{x}\sim\pi}\left[\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{2}+t^{2}}{2\left(\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{2}+t^{2}\right)^{3/2}}\right]$
@@
$\max_{\alpha}\mathbb{E}_{\mathbf{x}\sim\pi}\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{-1}\leq C_{\pi}.$
@@
$\mathbb{E}_{\mathbf{x}\sim\pi}\left[1-\frac{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}}{\sqrt{\left\|\mathbf{D}_{\alpha}\mathbf{x}\right\|_{2}^{2}+\varepsilon}}\right]\leq\frac{1}{2}C_{\pi}\varepsilon^{1/2}.$
Substitute this into Eq. 33, we have
$\left\|\nabla_{j}\log\pi(\mathbf{x})-\nabla_{j}\log\pi_{\varepsilon}(\mathbf{x})\right\|_{L^{1}(\pi)}\leq C_{\pi}\delta(m+2)^{2}\varepsilon^{1/2}.$
Finally, substitute this into Eq. 32, we have
@@
$2\lambda^{-1}c^{-1}\cdot C_{\pi}\delta(m+2)^{2}\varepsilon^{1/2}\leq C_{\pi}\cdot\frac{(m+2)^{2}}{32}\cdot\varepsilon=:C\varepsilon,$

### v3

@@ 6.1.3 Dimension-independent block acceptance rate
To test the
dimension-independent block acceptance rate in Proposition 4.1, we partition the original $512\times 512$ image into 4 sections of sizes $128\times 128$, $256\times 256$, $384\times 384$, and $512\times 512$. Furthermore, each section is partitioned into blocks of equal size $64\times 64$. Thus, the number of blocks in the sections of sizes $128\times 128$, $256\times 256$, $384\times 384$, and $512\times 512$ are $4$, $16$, $36$, and $64$, respectively. The 4 deblurring problems are shown on the left in Fig. 3.
We run the local & parallel MLwG given in Algorithm 2 with a step size of $\tau=7.44\times 10^{-6}$ on the 4 deblurring problems with different sizes. The step size is taken from a pilot run on the $512\times 512$ problem by targeting an acceptance rate of $0.547$ in each block and then taking the average of all block step sizes. For all problem sizes, we use a burn-in period of $31,250$ samples. We plot the acceptance rate for each block in Fig. 5 and see that the block acceptance rates are indeed dimension-independent.
@@ 6.1.4 Comparison to MALA
In this subsection, we compare the performance of our local & parallel MLwG (Algorithm 2) with MALA. For MALA, we use again the diminishing step size adaptation from [20] during burn-in, where we target an acceptance rate of $0.547$. The numbers of burn-in samples are listed in Table 2 and are chosen such that they increase linearly with the problem size. For MLwG, we use the same setting as in the previous tests.
We compare the sampling performance of MALA and MLwG in Table 2. In general, MLwG yields much larger nESS than MALA because it allows for a larger step size. In MALA, the proposal is global, and it is known that one should take $h=O(d^{-1/3})$ to ensure a non-degenerate acceptance rate [30]. In contrast, in MLwG, the proposal is local, and the step size can be taken independent of the total dimension, see Proposition 4.1. Furthermore, the nESS of MLwG becomes even larger as the problem size increases. We attribute this to the diminishing constraining effect of the boundary condition associated with the convolution operator on the inner blocks as the dimension increases. In addition, we note that for the given burn-in, MALA does not converge for the problem sizes $384\times 384$ and $512\times 512$, since the corresponding $\mathrm{max}~{}\mathrm{PSRF}>1.1$.
Table 2: Comparison of the local & parallel MLwG and MALA. The printed min nESS is the pixel-wise minimum of the mean nESS which is taken over the $5$ chains. For MLwG, the shown step size $\tau$ and acceptance rate $\alpha$ are the means taken over all blocks and the $5$ chains. The maximum PSRF is with respect to the pixels.
@@
min nESS [%] MLwG 2.6 2.8 2.9 3.0
MALA 1.4 0.8 0.6 0.4
@@
Notice that the results from Table 2 also validate the dimension-independent convergence rate in Proposition 4.2 of MLwG. This is because MLwG produces for all problem sizes and with the same burn-in converged chains with roughly constant PSRF. In contrast, MALA requires significantly more burn-in with increasing dimension. We note that despite the clear differences in sampling performance between MLwG and MALA, we do not see clear differences in the sample mean or the widths of the $90\%$ CIs. Therefore, we do not include these results here.
@@
We show the computing times in seconds per $1000$ samples in Fig. 6 and observe that the wall-clock time of MLwG remains almost constant and does not increase with the problem dimension. This is because the main computational effort of updating the $64\times 64$ blocks on each core remains constant and only more time is required for handling the increasing number of cores by the main process. For small problem sizes, the wall-clock time of MLwG is longer than that of MALA, because of the overhead of the parallelized implementation and the additional convolutions of fixed pixels in the local block likelihoods. However, since several updates are run in parallel in MLwG, its wall-clock time is eventually shorter than that of MALA, see the time for problem size $512\times 512$. Note that the total wall-clock time of MALA is actually significantly larger, since it requires much more burn-in. The benefits of MLwG obviously come at the cost of CPU time, which increases linearly with the number of cores.
@@ 6.2 House
In this section, we use another test image with a different blurring kernel to compare our local & parallel MLwG (Algorithm 2) with MALA.
@@ 6.2.1 Problem description
The degraded image is synthetically obtained via the observation model Eq. 3, where $\bm{A}\in\mathbb{R}^{d\times d}$ is obtained through the discrete PSF of a motion blur kernel with the length $17$ and the angle $45^{\circ}$. The noise is a realization of $\bm{\epsilon}\sim\mathcal{N}(\bm{0},0.01^{2}\bm{I})$. The true and the degraded image of “house” with the size $512\times 512$ are shown in the first row of Fig. 7.
Figure 7: The true image, the degraded image, the posterior sample mean, and the width of the 90% sample CI.
We use the adaptive total variation approach in [24] to determine the rate parameter $\delta$ in the TV prior Eq. 7 and obtain $\delta=34.83$.
@@ 6.2.2 Posterior sampling via MLwG and MALA
Again, we apply both the local & parallel MLwG and MALA to sample the smoothed posterior defined in Eq. 14. In both methods, we target an acceptance rate of $0.547$ using again the method from [20], where we adapt the step size during burn-in. In MLwG, we adapt the step size individually in each block. In the second row of Figure 7, we show the posterior mean and the width of the 90% CI obtained via the samples from MLwG. The restoration results from MALA are neglected, since they are visually identical to the ones from MLwG. However, the data about the sample chains that we show in Table 3 reveals once again the superior performance of MLwG over MALA. Similar as observed in the previous test, MLwG allows for a significantly larger step size, which leads to faster convergence and less correlated samples in terms of nESS.
Table 3: Sampling data of MLwG and MALA for the “house” image. The printed min nESS is the pixel-wise minimum of the mean nESS which is taken over the $5$ chains. For MLwG, the shown step size $\tau$ and acceptance rate $\alpha$ are the means taken over all blocks and the $5$ chains. The maximum PSRF is with respect to the pixels.
min nESS [%] $\bm{\tau}$ [10-6] $\bm{\alpha}$ [%] burn-in [103] max PSRF
MLwG 4.5 7.4 54.3 31.250 1.02
MALA 0.8 1.4 55.9 2000.000 1.09
@@ 7 Conclusions
Uncertainty quantification in imaging problems is usually a difficult task due to the high dimensionality of images. For image deblurring with TV prior, we presented a dimension-independent MLwG sampling algorithm. By exploiting the sparse conditional structure of the posterior, the proposed algorithm has dimension-independent block acceptance and convergence rates, which have been theoretically proven and numerically validated. To enable the use of the MALA proposal density, we used a smooth approximation of the TV prior. We showed that the introduced error is uniformly distributed over the pixels and thus is dimension-independent. Moreover, through numerical studies, we found that the smoothed posterior converges quickly to the exact posterior and yields feasible results for uncertainty quantification.
Various extensions of the local MALA-within-Gibbs method can be explored for imaging problems. For instance, the TV prior can be generalized to other difference operator-based priors that are of practical interest. On the implementation side, the local and parallel sampling nature of the method enables efficient GPU acceleration. We leave these considerations for future work.
@@ Acknowledgments
The work of RF and YD has been supported by a Villum Investigator grant (no. 25893) from The Villum Foundation. The work of SL has been partially funded by Singapore MOE grant A-8000459-00-00. The work of XT has been funded by Singapore MOE grants A-0004263-00-00 and A-8000459-00-00. We thank the anonymous reviewers for their insightful questions and comments, which contributed to a significant improvement of the manuscript.
@@ A.1 Dimension-independent approximation error
Due to the Kantorovich duality, the
Wasserstein-$1$ distance of $\pi_{i}$ and $\pi_{\varepsilon,i}$ can be written as
$W_{1}(\pi_{i},\pi_{\varepsilon,i})=\sup_{\phi_{0}\in{\rm Lip}_{1}(\mathbb{R}^{q})}\int\phi_{0}({\bm{x}}_{i})\left(\pi_{i}({\bm{x}}_{i})-\pi_{\varepsilon,i}({\bm{x}}_{i})\right)\mathrm{d}{\bm{x}}_{i},$
@@
$\phi({\bm{x}}_{i})=\phi_{0}({\bm{x}}_{i})-\int\phi_{0}({\bm{x}}_{i})\pi_{\varepsilon,i}({\bm{x}}_{i})\mathrm{d}{\bm{x}}_{i}=\phi_{0}({\bm{x}}_{i})-\int\phi_{0}({\bm{x}}_{i})\pi_{\varepsilon}({\bm{x}})\mathrm{d}{\bm{x}}.$
Then, $\phi\in{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})$ is the class of $1$-Lipschitz functions that are mean-zero w.r.t. $\pi_{\varepsilon,i}$:
(30) ${\rm Lip}_{1}^{0}(\pi_{\varepsilon,i}):=\left\{\phi\in{\rm Lip}_{1}(\mathbb{R}^{q}):\int\phi({\bm{x}}_{i})\pi_{\varepsilon,i}({\bm{x}}_{i})\mathrm{d}{\bm{x}}_{i}=0\right\}.$
Then, by definition,
$\int\phi_{0}({\bm{x}}_{i})\left(\pi_{i}({\bm{x}}_{i})-\pi_{\varepsilon,i}({\bm{x}}_{i})\right)\mathrm{d}{\bm{x}}_{i}=~{}$ $\int\phi({\bm{x}}_{i})\left(\pi_{i}({\bm{x}}_{i})-\pi_{\varepsilon,i}({\bm{x}}_{i})\right)\mathrm{d}{\bm{x}}_{i}$
@@
Given $\phi({\bm{x}}_{i})$, consider the Poisson equation for $u({\bm{x}})$:
(31) $\Delta u({\bm{x}})+\nabla\log\pi_{\varepsilon}({\bm{x}})\cdot\nabla u({\bm{x}})=\phi({\bm{x}}_{i}).$
@@
$\int\phi({\bm{x}}_{i})\pi({\bm{x}})\mathrm{d}{\bm{x}}=~{}$ $\int\left(-\nabla u({\bm{x}})\cdot\nabla\log\pi({\bm{x}})+\nabla\log\pi_{\varepsilon}({\bm{x}})\cdot\nabla u({\bm{x}})\right)\pi({\bm{x}})\mathrm{d}{\bm{x}}$
@@
Combining the above two inequalities, we have
(32) $\begin{split}\max_{i}W_{1}(\pi_{i},\pi_{\varepsilon,i})=~{}&\max_{i}\sup_{\phi\in{\rm Lip}_{1}^{0}(\pi_{\varepsilon,i})}\int\phi({\bm{x}}_{i})\pi({\bm{x}})\mathrm{d}{\bm{x}}\\ \leq~{}&2\lambda^{-1}c^{-1}\max_{j}\left\|\nabla_{j}\log\pi_{\varepsilon}-\nabla_{j}\log\pi\right\|_{L^{1}(\pi)}.\end{split}$
Note here, $\nabla_{j}\log\pi$ is not defined pointwise, but it suffices to require that $\nabla_{j}\log\pi\in L^{1}(\pi)$. It suffices to control the right hand side of Eq. 32. By definition,
$\nabla_{j}\log\pi({\bm{x}})-\nabla_{j}\log\pi_{\varepsilon}({\bm{x}})=\nabla_{j}\varphi_{\varepsilon}({\bm{x}})-\nabla_{j}\varphi_{0}({\bm{x}})$
@@
$\qquad\cdot\left(\left({\bm{D}}_{\alpha}^{(v)}{\bm{x}}\right)\left({\bm{D}}_{\alpha,j}^{(v)}\right)^{\rm T}+\left({\bm{D}}_{\alpha}^{(h)}{\bm{x}}\right)\left({\bm{D}}_{\alpha,j}^{(h)}\right)^{\rm T}\right)=:\delta\sum_{\alpha=1}^{d}{\bm{v}}_{\alpha}({\bm{x}}).$
Here, we denote ${\bm{D}}_{\alpha,j}^{(v)}\in\mathbb{R}^{1\times q}$ as the $j$-th block of ${\bm{D}}_{\alpha}\in\mathbb{R}^{1\times d}$ and similar for ${\bm{D}}_{\alpha,j}^{(h)}$. To control $\left\|{\bm{v}}_{\alpha}({\bm{x}})\right\|_{2}$, notice
$\left\|\left({\bm{D}}_{\alpha}^{(v)}{\bm{x}}\right)\left({\bm{D}}_{\alpha,j}^{(v)}\right)^{\rm T}+\left({\bm{D}}_{\alpha}^{(h)}{\bm{x}}\right)\left({\bm{D}}_{\alpha,j}^{(h)}\right)^{\rm T}\right\|_{2}\leq|{\bm{D}}_{\alpha}^{(v)}{\bm{x}}|\|{\bm{D}}_{\alpha,j}^{(v)}\|_{2}+|{\bm{D}}_{\alpha}^{(h)}{\bm{x}}|\|{\bm{D}}_{\alpha,j}^{(h)}\|_{2}$
$\leq\sqrt{2}\left(|{\bm{D}}_{\alpha}^{(v)}{\bm{x}}|+|{\bm{D}}_{\alpha}^{(h)}{\bm{x}}|\right)\leq 2\left(({\bm{D}}_{\alpha}^{(v)}{\bm{x}})^{2}+({\bm{D}}_{\alpha}^{(h)}{\bm{x}})^{2}\right)^{1/2}=:2\left\|{\bm{D}}_{\alpha}{\bm{x}}\right\|_{2}$
Here, we denote ${\bm{D}}_{\alpha}{\bm{x}}=({\bm{D}}_{\alpha}^{(v)}{\bm{x}},{\bm{D}}_{\alpha}^{(h)}{\bm{x}})\in\mathbb{R}^{2}$. Notice ${\bm{D}}_{\alpha,j}^{(v)}={\bm{D}}_{\alpha,j}^{(h)}=0$ if $\alpha$ does not live in block $j$ or a neighbor of block $j$, so that there are at most $(m+2)^{2}$ indices $\alpha$ s.t. ${\bm{v}}_{\alpha}({\bm{x}})$ is nonzero. Therefore,
(33) $\begin{split}\|\nabla_{j}\log\pi({\bm{x}})-\nabla_{j}\log&\pi_{\varepsilon}({\bm{x}})\|_{L^{1}(\pi)}\leq\delta\cdot\mathbb{E}_{{\bm{x}}\sim\pi}\sum_{\alpha:{\bm{v}}_{\alpha}({\bm{x}})\neq 0}\left\|{\bm{v}}_{\alpha}({\bm{x}})\right\|_{2}\\ \leq~{}&2\delta(m+2)^{2}\max_{\alpha}\mathbb{E}_{{\bm{x}}\sim\pi}\left[1-\frac{\left\|{\bm{D}}_{\alpha}{\bm{x}}\right\|_{2}}{\sqrt{\left\|{\bm{D}}_{\alpha}{\bm{x}}\right\|_{2}^{2}+\varepsilon}}\right].\end{split}$
Denote the function $I_{\alpha}(t):=\mathbb{E}_{{\bm{x}}\sim\pi}\left[1-\frac{\left\|{\bm{D}}_{\alpha}{\bm{x}}\right\|_{2}}{\sqrt{\left\|{\bm{D}}_{\alpha}{\bm{x}}\right\|_{2}^{2}+t^{2}}}\right]$. Then $I_{\alpha}(0)=0$ and
$I_{\alpha}^{\prime}(t)=~{}$ $\mathbb{E}_{{\bm{x}}\sim\pi}\left[\frac{\left\|{\bm{D}}_{\alpha}{\bm{x}}\right\|_{2}t}{\left(\left\|{\bm{D}}_{\alpha}{\bm{x}}\right\|_{2}^{2}+t^{2}\right)^{3/2}}\right]\leq\mathbb{E}_{{\bm{x}}\sim\pi}\left[\frac{\left\|{\bm{D}}_{\alpha}{\bm{x}}\right\|_{2}^{2}+t^{2}}{2\left(\left\|{\bm{D}}_{\alpha}{\bm{x}}\right\|_{2}^{2}+t^{2}\right)^{3/2}}\right]$
@@
$\max_{\alpha}\mathbb{E}_{{\bm{x}}\sim\pi}\left\|{\bm{D}}_{\alpha}{\bm{x}}\right\|_{2}^{-1}\leq C_{\pi}.$
@@
$\mathbb{E}_{{\bm{x}}\sim\pi}\left[1-\frac{\left\|{\bm{D}}_{\alpha}{\bm{x}}\right\|_{2}}{\sqrt{\left\|{\bm{D}}_{\alpha}{\bm{x}}\right\|_{2}^{2}+\varepsilon}}\right]\leq\frac{1}{2}C_{\pi}\varepsilon^{1/2}.$
Substituting this into Eq. 33, we have
$\left\|\nabla_{j}\log\pi({\bm{x}})-\nabla_{j}\log\pi_{\varepsilon}({\bm{x}})\right\|_{L^{1}(\pi)}\leq C_{\pi}\delta(m+2)^{2}\varepsilon^{1/2}.$
Finally, substituting this into Eq. 32, we have
@@
$2\lambda^{-1}c^{-1}\cdot C_{\pi}\delta(m+2)^{2}\varepsilon^{1/2}\leq C_{\pi}\cdot\frac{(m+2)^{2}}{32}\cdot\varepsilon=:C\varepsilon$